# Real-Time Flying Object Detection with YOLOv8

### v1

@@ 3.2 Model Examples and Results
3.3 Refined Model - Transfer Learning
@@
After applying transfer learning to our “real world“ data set [1], the results indicate that our model is a solid foundation for transfer learning and effectively extracted flying object feature representations. This data set was selected to focus on our challenge of detecting and classifying extremely small objects in appearance. Figure LABEL:fig:RefinedModel displays our results, featuring four distinct images that represent the bird, drone, airplane, and helicopter objects.
@@
The first image displays a small, pixelated bird that only takes up 0.02% of the image. Even with the lack of the bird class in our training process, our model correctly identified that the object was not any of the other available classes, even while allowing a very low confidence threshold of 0.20. The second image contains a drone, which also only took up 0.02% of its image. This drone is nearly indistinguishable from the background clouds to the human eye, yet our model was still able to classify it with a confidence score of 0.81. The third image includes a small airplane that takes up 0.034% of pixels, which our model was still able to correctly identify and classify with a high confidence score of 0.85. In the final image, a barely visible helicopter (0.01% of the image) was correctly classified with a confidence score of 0.73.
@@
Figure 3 we can see that in the first layer, the feature map activation correctly segments the object. The second layer starts picking out all of the tree tops which can be explained by the higher relative variance of the tree tops. In the third layer, we see mroe importance being placed on the background and more granular features being detected. In the fourth layer, we see the outline of the drone itself.
Moving on to the second row, we can see the true strength of the localization accuracy with an over-emphasized detection. In the second layer, we see a de-emphasis on the background. In the third and fourth layer, we see the same behavior as before.
@@
Our model performed exceptionally well, even while presented with the challenges of size, varying flying objects, and camouflaged objects that would be difficult for the human eye to identify. These results demonstrate that our model serves as an excellent base for transfer learning, particularly when dealing with small, pixelated objects, blended backgrounds, and distinguishing between drones and other flying objects. After 190 epochs with the weights learned from the generalized model as the initialization, the refined model achieves an mAP50-95 of 0.835 across the plane, helicopter, and drone classes.
@@ 4 Model Architecture
The main proposal from YOLO is to frame object detection as a one-pass regression problem. YOLOv1 comprises a single neural network, predicting bounding boxes and associated class probability in a single evaluation. The base model of YOLO works by first dividing the input image into an S x S grid where each grid cell (i,j) predicts B bounding boxes, a confidence score for each box and C class probabilities. The final output will be a tensor of shape: S x S x (B x 5 + C).
@@ 4.1 YOLOv1 Overview
YOLOv1 architecture 5 consists of 24 convolutional layers followed by two fully connected layers. In the paper, the authors took the first 20 convolutional layers from the backbone of the network and, with the addition of an average pooling layer and a single fully connected layer, where it was pre-trained and validated on the ImageNet 2012 dataset. During inference, the final four layers and 2 FC layers are added to the network; all initialized randomly.
@@
YOLOv1 uses stochastic gradient descent as its optimizer; the Loss function is shown here 5. The Loss function 5 comprises two parts, localization loss, and classification loss. The localization loss measures the error between the predicted bounding box coordinates and the ground-truth bounding box. The classification loss measures the error between the predicted class probabilities and the ground truth. The $\lambda_{coord}$ and $\lambda_{noobj}$ are regularization coefficients that regulate the magnitude of the different components, emphasizing object localization and deemphasizing grid cells without objects.
@@ 4.2 YOLOv5 Overview
YOLOv5 [4] is an object detection model introduced in 2020 by Ultralytics, the originators of the original YOLOv1 and YOLOv3. YOLOv5 achieves SOTA performance on the COCO benchmark dataset [3] while also being fast and efficient to train and deploy. YOLOv5 made several architectural changes, most notably the standardized practice of structuring the model into three components, the backbone, neck, and head.
The backbone of YOLOv5 is Darknet53, a new network architecture that focuses on feature extraction characterized by small filter windows and residual connections. Cross Stage Partial connections enable the architecture to achieve a richer gradient flow while reducing computation as described [20] proposed by Wang et al.
@@
YOLOv5’s head consists of three branches, each predicting a different feature scale. In the original publication of the model [3], the creators used three grid cell sizes of 13 x 13, 26 x 26, and 52 x 52, which each grid cell predicting B = 3 bounding boxes. Each head produces bounding boxes, class probabilities, and confidence scores. Finally, the network uses Non-maximum Suppression (NMS) [9] to filter out overlapping bounding boxes.
YOLOv5 incorporates anchor boxes, fixed-sized bounding boxes used to predict the location and size of objects within an image. Instead of predicting arbitrary bounding boxes for each object instance, the model predicts the coordinates of the anchor boxes with predefined aspect ratios and scales and adjusts them to fit the object instance.
@@ 4.3 YOLOv8 Overview
YOLOv8 is the latest version of the YOLO object detection model. This latest version has the same architecture as its predecessors 6 but it introduces numerous improvements compared to the earlier versions of YOLO such as a new neural network architecture that utilizes both Feature Pyramid Network (FPN) and Path Aggregation Network (PAN) and a new labeling tool that simplifies the annotation process. This labeling tool contains several useful features like auto labeling, labeling shortcuts, and customizable hotkeys. The combination of these features makes it easier to annotate images for training the model.
The FPN works by gradually reducing the spatial resolution of the input image while increasing the number of feature channels. This results in the creation of feature maps that are capable of detecting objects at different scales and resolutions. The PAN architecture, on the other hand, aggregates features from different levels of the network through skip connections. By doing so, the network can better capture features at multiple scales and resolutions, which is crucial for accurately detecting objects of different sizes and shapes.  [18]
@@ 4.4 YOLOv8 vs YOLOv5
The reason YOLOv8 is being compared to YOLOv5 and not any other version of YOLO is that YOLOv5’s performance and metrics are closer to YOLOv8’s. However, YOLOv8 surpasses YOLOv5 in aspects including a better mAP as seen in Figure LABEL:fig:YOLOv8_mAP. Along with a better mAP, this shows that YOLOv8 has fewer outliers when measured against the RF100 which is a 100-sample dataset from the Roboflow universe which is a repository of 100,000 datasets. We also witness YOLOv8 outperforming YOLOv5 for each RF100 category. From Figure LABEL:fig:YOLOv8_average_mAP_against_cats we can see that YOLOv8 produces similar or even better results compared to YOLOv5  [16].
@@
As mentioned previously, YOLOv8 uses a new architecture that combines both FAN and PAN modules. FPN is used to generate feature maps at multiple scales and resolutions, while PAN is used to aggregate features from different levels of the network to improve accuracy. The results of the combined FAN and PAN modules are better than YOLOv5 which uses a modified version of CSPDarknet architecture. This modified version of CSPDarknet is based on the cross-stage partial connections (CSP), which improves the flow of information between different parts of the network.
Figure 7: YOLOv8 vs Previous Versions
@@
Another difference the two models have is their training data. YOLOv8 was trained on a larger and more diverse dataset compared to YOLOv5. YOLOv8 was trained on a blend of the COCO dataset and several other datasets, while YOLOv5 was trained primarily on the COCO dataset. Because of that, YOLOv8 has a better performance on a wider range of images.
@@
YOLOv8 includes a new labeling tool called RoboFlow Annotate which is used for image annotation and object detection tasks in computer vision. RoboFlow Annotate makes it easier to annotate images for training the model and includes several features such as auto labeling, labeling shortcuts, and customizable hotkeys. In contrast, YOLOv5 uses a different labeling tool called LabelImg. LabelImg is an open-source graphical image annotation tool that allows its users to draw bounding boxes around objects of interest in an image, and then export the annotations in the YOLO format for training the model.
@@
YOLOv8 includes more advanced post-processing techniques than YOLOv5, which is a set of algorithms applied to the predicted bounding boxes and objectiveness scores generated by the neural network. These techniques help to refine the detection results, remove redundant detections, and improve the overall accuracy of the predictions. YOLOv8 uses Soft-NMS which is a variant of the NMS technique used in YOLOv5. Soft-NMS applies a soft threshold to the overlapping bounding boxes instead of discarding them outright. Whereas NMS removes the overlapping bounding boxes and keeps only the ones with the highest objectiveness score.
@@
Output heads refer to the final layers of a neural network that predict the locations and classes of objects in an image. In YOLO architecture there are typically several output heads that are responsible for predicting different aspects of the detected objects, such as the bounding box coordinates, class probabilities, and objectiveness scores. These output heads are typically connected to the last few layers of the neural network and are trained to output a set of values that can be used to localize and classify objects in an image. The number and type of output heads used can vary depending on the specific object detection algorithm and the requirements of the task at hand. YOLOv5 has 3 output heads while YOLOv8 has 1 output head. YOLOv8 Does not have small, medium, and large anchor boxes rather it uses an anchor free detection mechanism that directly predicts the center of an object instead of the offset from a known anchor box which reduces the number of box predictions, and that speeds up the post processing process.
@@
It is fair to note that YOLOv8 is slightly slower than YOLOv5 in regards to object detection speed. However, YOLOv8 is still able to process images in real-time on modern GPUs.

### v2

@@ 3.1 Generalized Model
3.2 Refined Model
@@
To generate the refined model, we initialized the model with the weights learned from the generalized model and default hyperparameters. We then trained the model on the “real world“ data set for 199 epochs [1]. This data set was selected to focus on our challenge of detecting and classifying extremely small objects in appearance. Figure 5 displays our results, featuring four distinct images that represent the bird, drone, airplane, and helicopter objects.
@@
The first image contains an extremely small bird that only takes up 0.02% of the image. Even with the lack of the bird class in our training process, our model correctly identified that the object was not any of the other available classes, even while allowing a very low confidence threshold of 0.20. The second image contains a drone, which also only took up 0.02% of its image. This drone is nearly indistinguishable from the background clouds to the human eye, yet our model was still able to classify it with a confidence score of 0.81. The third image includes a small airplane that takes up 0.034% of pixels, which our model was still able to correctly identify and classify with a high confidence score of 0.85. In the final image, a barely visible helicopter (0.01% of the image) was correctly classified with a confidence score of 0.73.
@@
In Figure 4, we can see that the feature map activation correctly segments the object in the first layer. The second layer starts picking out all of the tree tops, which can be explained by the higher relative variance of the tree tops. In the third layer, we see more importance being placed on the background and more granular features being detected. In the fourth layer, we see the outline of the drone itself. In the second row, the true strength of the localization accuracy with an over-emphasized detection is displayed. In the second layer, we see a de-emphasis on the background. In the third and fourth layers, we see the same behavior as before.
@@
Our model achieves state-of-the-art results, achieving a mAP50 of 0.991 and mAP50-95 of 0.835 across the plane, helicopter, and drone classes. These results demonstrate that our generalized model serves as an excellent base for transfer learning, particularly when dealing with extremely small objects, blended backgrounds, and distinguishing between drones and other flying objects.
4 Discussion
To the problem of flying object detection, we apply transfer learning with weights learned from our generalized model to our refined model in order to achieve state-of-the-art results in this domain. We argue that our algorithm extracts better feature representations of flying objects than those seen in previous research, furthering the current state of research in this domain. Our refined model achieves a 99.1% mAP50, 98.7% Precision, and 98.8% Recall with 50 fps inference speed on the 3-class data set (drone, plane, and helicopter), surpassing models generated from previous research to a significant extent. Aydin et al. proposed a YOLOv5 instance that achieved 90.40% mAP50, 91.8% Precision, and 87.5% Recall with 31 fps inference speed trained on a data set only containing drones and birds [3]. Rozantsev et al. trained their proposed model on a data set reflective of ours, containing flying objects that occupy small portions of the input image with clustered backgrounds. They achieve an 84.9% AP on a data set containing only UAVs and 86.5% AP on a data set containing only aircraft [18]. Al-Qubaydhi et al. proposed a model utilizing the YOLOv5 framework and achieves an impressive 94.1% mAP50, 94.7% Precision, and 92.5% Recall on a dataset containing only one class of drones. [2]. Even with our exceptional results, a potential limitation of our refined model is that it was trained on a data set with a low amount of distinct environments. To address this potential generalization issue, we suggest utilizing our generalized model weights to transfer learn on a data set with higher frequency of distinct backgrounds.
@@ 5 Model Architecture
The main proposal from YOLO is to frame object detection as a one-pass regression problem. YOLOv1 comprises a single neural network, predicting bounding boxes and associated class probability in a single evaluation. The base model of YOLO works by first dividing the input image into an S x S grid where each grid cell (i,j) predicts B bounding boxes, a confidence score for each box, and C class probabilities. The final output will be a tensor of shape S x S x (B x 5 + C).
@@ 5.1 YOLOv1 Overview
YOLOv1 architecture [Figure 6] consists of 24 convolutional layers followed by two fully connected layers. In the paper [14], the authors took the first 20 convolutional layers from the backbone of the network and, with the addition of an average pooling layer and a single fully connected layer, it was pre-trained and validated on the ImageNet 2012 dataset. During inference, the final four layers and 2 FC layers are added to the network; all initialized randomly.
@@
YOLOv1 uses stochastic gradient descent as its optimizer. The loss function, shown by Equation 5, comprises two parts: localization loss and classification loss. The localization loss measures the error between the predicted bounding box coordinates and the ground-truth bounding box. The classification loss measures the error between the predicted class probabilities and the ground truth. The $\lambda_{coord}$ and $\lambda_{noobj}$ are regularization coefficients that regulate the magnitude of the different components, emphasizing object localization and de-emphasizing grid cells without objects.
@@ 5.2 YOLOv5 Overview
YOLOv5 [6] is an object detection model introduced in 2020 by Ultralytics, the originators of the original YOLOv1 and YOLOv3. YOLOv5 achieves state-of-the-art performance on the COCO benchmark dataset [5] while also being fast and efficient to train and deploy. YOLOv5 made several architectural changes, most notably the standardized practice of structuring the model into three components: the backbone, neck, and head.
The backbone of YOLOv5 is Darknet53, a new network architecture that focuses on feature extraction characterized by small filter windows and residual connections. Cross-stage partial connections (CSP) enable the architecture to achieve a richer gradient flow while reducing computation as proposed by Wang et al. [23].
@@
YOLOv5’s head consists of three branches, each predicting a different feature scale. In the original publication of the model [5], the creators used three grid cell sizes of 13 x 13, 26 x 26, and 52 x 52, with each grid cell predicting B = 3 bounding boxes. Each head produces bounding boxes, class probabilities, and confidence scores. Finally, the network uses Non-maximum Suppression (NMS) [11] to filter out overlapping bounding boxes.
YOLOv5 incorporates anchor boxes, which are fixed-sized bounding boxes used to predict the location and size of objects within an image. Instead of predicting arbitrary bounding boxes for each object instance, the model predicts the coordinates of the anchor boxes with predefined aspect ratios and scales and adjusts them to fit the object instance.
@@ 5.3 YOLOv8 Overview
YOLOv8 is the latest version of the YOLO object detection model. This latest version has the same architecture as its predecessors [Figure 7], but it introduces numerous improvements compared to the earlier versions of YOLO, such as a new neural network architecture that utilizes both Feature Pyramid Network (FPN) and Path Aggregation Network (PAN) and a new labeling tool that simplifies the annotation process. This labeling tool contains several useful features like auto labeling, labeling shortcuts, and customizable hotkeys. The combination of these features makes it easier to annotate images for training the model.
The FPN works by gradually reducing the spatial resolution of the input image while increasing the number of feature channels. This results in the creation of feature maps that are capable of detecting objects at different scales and resolutions. The PAN architecture, on the other hand, aggregates features from different levels of the network through skip connections. By doing so, the network can better capture features at multiple scales and resolutions, which is crucial for accurately detecting objects of different sizes and shapes  [21].
@@ 5.4 YOLOv8 vs YOLOv5
The reason why YOLOv8 is being compared to YOLOv5 rather than any other version of YOLO is that YOLOv5’s performance and metrics are closer to YOLOv8’s. However, YOLOv8 surpasses YOLOv5 in aspects including a better mAP as seen in Figure 9. Along with a better mAP, this shows that YOLOv8 has fewer outliers when measured against the RF100. RF100 is a 100-sample dataset from the Roboflow universe, which is a repository of 100,000 data sets. We also witness YOLOv8 outperforming YOLOv5 for each RF100 category. From Figure 9, we can see that YOLOv8 produces similar or even better results compared to YOLOv5  [19].
@@
As mentioned previously, YOLOv8 uses a new architecture that combines both FAN and PAN modules. FPN is used to generate feature maps at multiple scales and resolutions, while PAN is used to aggregate features from different levels of the network to improve accuracy. The results of the combined FAN and PAN modules are better than YOLOv5, which uses a modified version of the CSPDarknet architecture. This modified version of CSPDarknet is based on cross-stage partial connections, which improves the flow of information between different parts of the network.
Figure 8: YOLOs mAP@.50 against RF100.
Figure 9: YOLOs average mAP@.50 against RF100 categories
@@
Another difference the two models have is based on their training data. YOLOv8 was trained on a larger and more diverse dataset compared to YOLOv5. YOLOv8 was trained on a blend of the COCO dataset and several other datasets, while YOLOv5 was trained primarily on the COCO dataset. Because of this, YOLOv8 has a better performance on a wider range of images.
@@
YOLOv8 includes a new labeling tool called RoboFlow Annotate, which is used for image annotation and object detection tasks in computer vision. RoboFlow Annotate makes it easier to annotate images for training the model and includes several features such as auto labeling, labeling shortcuts, and customizable hotkeys. In contrast, YOLOv5 uses a different labeling tool called LabelImg. LabelImg is an open-source graphical image annotation tool that allows its users to draw bounding boxes around objects of interest in an image, and then export the annotations in the YOLO format for training the model.
@@
YOLOv8 includes more advanced post-processing techniques than YOLOv5, which is a set of algorithms applied to the predicted bounding boxes and objectiveness scores generated by the neural network. These techniques help to refine the detection results, remove redundant detections, and improve the overall accuracy of the predictions. YOLOv8 uses Soft-NMS, a variant of the NMS technique used in YOLOv5. Soft-NMS applies a soft threshold to the overlapping bounding boxes instead of discarding them outright, whereas NMS removes the overlapping bounding boxes and keeps only the ones with the highest objectiveness score.
@@
Output heads refer to the final layers of a neural network that predict the locations and classes of objects in an image. In YOLO architecture, there are normally several output heads that are responsible for predicting different aspects of the detected objects, such as the bounding box coordinates, class probabilities, and objectiveness scores. These output heads are typically connected to the last few layers of the neural network and are trained to output a set of values that can be used to localize and classify objects in an image. The number and type of output heads used can vary depending on the specific object detection algorithm and the requirements of the task at hand. YOLOv5 has three output heads while YOLOv8 has one output head. YOLOv8 does not have small, medium, and large anchor boxes. It uses an anchor-free detection mechanism that directly predicts the center of an object instead of the offset from a known anchor box. This reduces the number of box predictions and speeds up the post-processing process in return.
@@
It is fair to note that YOLOv8 is slightly slower than YOLOv5 in regard to object detection speed. However, YOLOv8 is still able to process images in real-time on modern GPUs.